You have 15 coins in your pocket. The coins are either quarters or dimes.You have a total of $3.15. How many quarters and dimes do you have?
Found 2 solutions by ichigo449, richwmiller:
Answer by ichigo449(30)

Let x be the number of quarters and y dimes. Then x+y = 15 and, in units of pennies, 25x+10y = 315. Now substitute y int terms of x: 25x+10(15-x) = 315 or 25x-10x+150 = 315 or 15x = 165 so x = 11. Now y must be 4 so there are 11 quarters and 4 pennies. Good day.

Answer by richwmiller(17219)

q+d=15
q=15-d
25q+10d=315
25(15-d)+10d=315
375-25d+10d=315
-15d=-60
dimes =4
quarters=11